Q3[2 marks]: Assume that youre the receiver and that the 4B/5B block coding is used. Youre given all the mapping between 4
Answer #1

Answer of Q3.

Image is blur so i can't understand

Original bits-01001000101101011101

Answer of Q4.

Given:-BW=12KHz , mid point=29KHz

So,

Lower Frequency = mid point - BW/2

=29 - 12/2

= 29 - 6

=23KHz

Higher Frequency=mid point + BW/2

=29 + 12/2

= 29 + 6

= 35KHz